### v1

@@ 4 Conclusion
This work investigates on prognosis prediction of cancer patients using portrait photos. Our work demonstrates that the latent features extracted by StyleGAN contain highly relevant prognostic information, and an effective regression model can be optimized for survival prediction. Such a model using facial features have good interpretability and can achieve comparable survival prediction accuracy to a state-of-art model using clinical record data. The constructed health attribute allows photo editing into healthier appearance or prognosis of a more severe illness, which has important potential value for patient care.

### v3

@@ 4 Conclusion
This work investigates on prognosis prediction of cancer patients using portrait photos. Our work demonstrates that the latent features extracted by StyleGAN contain highly relevant prognostic information, and an effective regression model can be optimized for survival prediction. Such a model using facial features has good interpretability and can achieve comparable survival prediction accuracy to a state-of-the-art model using clinical record data. The constructed health attribute allows photo editing into a healthier appearance or prognosis of a more severe illness, which has important potential value for patient care.
[…]
The authors have no conflict of interest to declare for this work.